suppose the consumer wants to consume 2 goods . Price of goods are rs 4 and rs5 respectively . The consumers income is rs 20

write down the equation for budget line



1
Expert's answer
2021-02-02T15:40:10-0500

Answer

The consumer has a given income which sets limits to his maximum behavior.

The equation of the lineis

P1x1+p2x2=M

According to the question

P1=4, p2=5

4x+5y=20

Budget set may be

Two combination possible

1) (5,0)

2) (0,4)

4*5+5*0=20

4*0+5*4=20

Another way

The quantity of goodx purchased if whole of the given income 20 is spent on it is ox. Therefore

OX*px= 20

"OX=20\/4 =5"

Now, the quantity of goody purchased if whole of the given income 20 is spent on it is OY. Therefore,

OY*P = M OR

"OY= 20\/5 =4"

Dividing (2) by (1) we have:

"OY\/OX =M\/P1\/M\/P2"


"= 20\/5*4\/20"

"4\/5"


"(-) Px\/Py"

It is thus proved that the slope of the price line BL represents the ratio of the price of two goods.

Budget line equation

"PxQx+Py Qy =Money income"